OpenStudy (anonymous):

-11x^2 = x + 11 what is the number of real solutions

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

howdy

OpenStudy (anonymous):

\[-11x^2 = x + 11\] \[11x^2+x+11=0\] you can check the number of real solutions by computing \[b^2-4ac\] which in this case is \(1^2-4\times 11\times 11\)

OpenStudy (anonymous):

since the answer is clearly negative there are no real solutions